Find the value of σ x̅. Use the choices on no. 4 problem. Use this data: Consider the score examples illustrated, in which random samples of size 16 are obtained from N (25,9)


1
Expert's answer
2022-04-13T09:33:54-0400

"\sigma_{\bar X}=\frac{\sigma}{\sqrt{n}}=\frac{3}{\sqrt{16}}=0.75."